I need to figure out how many pounds of Dry Ice or liters of Liquid Nitrogen are needed to cool 1 cubic inch of both Aluminum from room temp 23 deg C to -70 deg C.
1
Expert's answer
2017-04-11T01:51:07-0400
The heat required to cool down an inch of Al is defined as Q = mC(t1 - t2) = 2700 kg/m3 * 1.6*10^-5 m3 * 903 J/kg/K * 93 K = 3.6 kJ.
Thus, the heat of vaporization for liq. N2 is 198 kJ/kg and to cool down the piece of Al we need 3.6 kJ/198 kJ/kg = 0.018 kg of it.
In pounds it is some 0.018 kg/0.454 kg/pound = 0.04 pound of liquid nitrogen.